A stone falls from a balloon that is descending at a uniform rate of 12m/s. the displacement from the stone from the point of release after 10 sec is

(a) 490 m

(b) 510m

(c) 610 m

(d) 725 m

Grade:9

6 Answers

jishnu prakash
13 years ago

The point of release is a point in the air, not attached to the balloon.

Since speed of balloon has same direction as acceleration due to gravity,

u = 12m/s    a=9.8 m2/s  t=10s

s = ut  +  0.5at2 

  = 120   +   (0.5)(9.8)(100)

  = 610 m

 

vikas askiitian expert
13 years ago

since the stone is falls from a stone which is decending with 12m/s

 that means unitial velocity of stone is given 12m/s .......

for constant accleration we have , s(displacement) = ut + at2/2

a = -g = -9.8m/s2(downward)      ,  u = -12m/s (downward)

             s = -12t - 9.8t2/2

after 10 sec

          s = -120-490=610m

option c is correct

 

divyanshu
8 years ago
610 m
Anubhav
6 years ago
Since balloon is descending at rate of 12m/s therefore the initial velocity of stone will be equal to 12m/sBy distance formulaS=ut+½at²=12*10+½*9.8*81=610
swetlana
6 years ago
we know that s=ut+1/2at2 
here, u=12m/s; a=9.8m/s2; t=10 s
by substituting the values, we get
s=12×10+1/2(9.8)(10)2
s=610 m
Therefore,the displacement of the stone from the point of release after 10 seconds is 610 metres.
helper
2 years ago
Distance(y),s=ut+1/2at^2
This is the formulae we would be working on...
And to solve this question we have just follow some easy steps.
 
So first now we will look at the values we have been given.
A stone falls from a balloon that is descending at a uniform rate of   12m/s   (u/initial velocity)  . the displacement from the stone from the point of release after  10 sec  (time/t!)    is...
So here goes out question. So we found out easily the time and u by just looking at the question. So we will just substitute the values and get a cool answer!!
 
s=ut+1/2at^2
 
(-s)=  (-12)*10  +   1/2*(-9.8)*(10)^2
 
Note. A negative sign means the action is in the downward direction.
 
So crossing out all the negative marks by taking them common we get
s=12*10  +   9.8*1/2*(10)^2
 
Solve it!
 
So finally we get that, s=610m
 
Great job! We did it! So, before moving on to the next question just a bit of kind advice-
 
 
“Read the question carefully! The answer is always hidden in the question itself”